determine the equation of the line passing through the points (4,-3) and (-4,6)
Answer by checkley71(8403)

FIRST WE FIND THE SLOPE BY (Y2-Y1)/(X2-X1)
(6+3)/(-4-4)
9/-8=-9/8 NOW SUBSTITUTING ANY SET OF X&Y VALUES INTO THE LINE EQUATION
Y=mX+b WE FIND THE Y INTERCEPT VALUE (b)
-3=-9/8*4+b
-3=-36/8+b
-3=-4.5+b
b=-3+4.5
b=1.5 THUS THE LINE EQUATION IS
Y=-9/8X+1.5